OpenStudy (anonymous):

I got this question wrong on my test. Could some one please show me the steps? Choose the equation below that represents the line passing through the point (-2,-3) with a slope of -6. Choices are y+3=-6(x+2) and y-3=-6(x-2).

OpenStudy (texaschic101):

y - y1 = m(x - x1) m = -6 y1 = -3 x1 = -2 now lets sub our values in y - (-3) = -6(x -(-2) now what can you do with this ?

OpenStudy (anonymous):

simplify to y-3=-6(x-2) it looks so easy written out like that. Thank you

OpenStudy (texaschic101):

wrong.... y -(-3) = -6(x - (-2) y + 3 = -6(x + 2) two negatives multiplied together equal a positive

OpenStudy (texaschic101):

understand ?

OpenStudy (anonymous):

I keep making the same mistake. There are no exceptions to the two neg rule?

OpenStudy (texaschic101):

keep practicing...you will get it. No exceptions...negative multiplied by negative = positive.

OpenStudy (anonymous):

ok thank you

OpenStudy (texaschic101):

anytime :)